2 moles of nitrogen is mixedvwith 6 mole of hydrogen in a closed vessrl if 50% nitrogen reacts to form nh3 then constant is

Grade:11

2 Answers

Arun
6 years ago
Y
Arun
6 years ago
N2 + 3H2 ----> 2NH3
Initial moles 2 2 0
Equilibrium moles 2-\alpha 2-3\alpha 2\alpha
Since 0.5 mole of N2 has been used up therefore :
\alpha= 0.5 mole
Moles of NH3 = 1
Concentration of NH3 = moles/volume = 1/5 mol/l